# Widths of embeddings of Gaussian Sobolev spaces

### 14 Apr 2026 v1

@@ 3 Approximation for Gaussian Sobolev spaces
$\int_{{\mathbb{I}}^{d}_{\theta}}\int_{{\mathbb{I}}^{d}_{\theta}}\frac{\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}}+{\boldsymbol{k}})-D^{\boldsymbol{\alpha}}f({\boldsymbol{y}}+{\boldsymbol{k}})\big|^{p}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\tilde{s}p}}{\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\ll\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\alpha}f({\boldsymbol{y}})\big|^{p}e^{\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{2}}{\rm d}\gamma({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{y}})$
@@
which implies
@@
The condition $q_{*}<\frac{p_{*}}{2}$ ensures that the sum is bounded by $Cn^{-a}(\log n)^{b}\|f\|_{W^{s}{p,G}(\mathbb{R}^{d},\gamma)}$. The remaining steps follow similarly to those in the proof for $W^{s}_{p}({\mathbb{R}}^{d},\gamma)$. The proof is complete.
@@
Proof. We present the proof for the case of Kolmogorov and linear widths; the case of sampling widths follows similarly. In this proof, for $k>0$ we denote ${\mathbb{U}}^{d}_{k}=[-k,k]^{d}$. Let $f$ be a $1$-periodic function on ${\mathbb{R}}^{d}$ and $f\in\tilde{W}^{\alpha}_{p}({\mathbb{I}}^{d})$. Denote by $f_{\operatorname{ext}}$ the extension of $f$ to $W^{s}_{p}({\mathbb{R}}^{d})$. For the existence of such a extension we refer the reader to [21, Theorem 1.105]. Without loss of generality we can assume that $\operatorname{supp}(f)\in{\mathbb{U}}^{d}_{1}$. Hence we have
@@
Hence we get
$\lambda_{n}\big({\boldsymbol{W}}^{s}_{p}(\mathbb{R}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\geq d_{n}\big({\boldsymbol{W}}^{s}_{p}(\mathbb{R}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\gg d_{n}(\tilde{{\boldsymbol{W}}}^{\alpha}_{p}({\mathbb{I}}^{d}),\tilde{L}_{q}({\mathbb{I}}^{d})).$
@@
    $d_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{I}}^{d}),L_{q}({\mathbb{I}}^{d})\big)\asymp\lambda_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{I}}^{d}),L_{q}({\mathbb{I}}^{d})\big)\asymp n^{-s/d}.$
@@
    $\varrho_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{I}}^{d}),L_{q}({\mathbb{I}}^{d})\big)\asymp n^{-s/d}.$
@@
$d_{n}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)=\lambda_{n}\big({\boldsymbol{W}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big).$
@@
where the linear operators $A$, $D_{\sqrt{\rho_{\alpha}}}$, and $B$ are defined as
@@
$D\xi$ $=(\xi_{\boldsymbol{k}}/(1+|{\boldsymbol{k}}|_{1})^{s/2})_{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}}\,,\qquad\xi=(\xi_{\boldsymbol{k}})_{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}}$
@@
$d_{n}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)\leq d_{n}\big(D\big)\,.$
@@
$d_{n}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)\geq d_{n}\big(D\big)\,.$
@@
$\lambda_{n}\big(Id\big)=\lambda_{n}\big(D\big).$
We know that $d_{n}\big(D\big)=a_{n}(D)=\sigma_{n}$ where $\sigma_{n}$ is the non-increasing rearrangement of $(1+|{\boldsymbol{k}}|_{1})^{-s/2}$. We know that the cardinality of the set $\{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}:1+|{\boldsymbol{k}}|_{1}\leq r\}=$ is $c(r,d)=\binom{r-1+d}{d}$. If $c(r-1,d)<n\leq c(r,d)$. Then we have
@@
Using $c(r,d)=\binom{r-1+d}{d}$ we obtain the second statement.
@@ 4 Gaussian Sobolev spaces with mixed smoothness
For $s\in{\mathbb{N}}$, we define the Gaussian Sobolev space $W^{s}_{p,{\rm m}}({\mathbb{R}}^{d},\gamma)$ as the normed space of all functions $f\in L_{p}({\mathbb{R}}^{d},\gamma)$ such that the generalized partial derivatives $D^{\boldsymbol{\alpha}}f$ of order $\boldsymbol{\alpha}$} belong to $L_{p}({\mathbb{R}}^{d},\gamma)$ for all ${\boldsymbol{\alpha}}\in{\mathbb{N}}_{0}^{d}$ satisfying $|{\boldsymbol{\alpha}}|_{\infty}\leq s$. The norm of a function $f$ in this space is defined by
@@
Let $1\leq p<\infty$, $s>0$ and $s\not\in{\mathbb{N}}$. Then the fractional Sobolev spaces of mixed smoothness $W_{p,{\rm m}}^{s}({\mathbb{U}}^{d})$ is defined by
@@
Let $1\leq p<\infty$, $s>0$ and $s\not\in{\mathbb{N}}$. We define the fractional Gaussian Sobolev space $W_{p,{\rm m}}^{s}(\mathbb{R}^{d},\gamma)$ of mixed smoothness as
$W_{p,{\rm m}}^{s}(\mathbb{R}^{d},\gamma):=\left\{f\in W^{\bar{s}}_{p,{\rm m}}(\mathbb{R}^{d},\gamma)\;;\;[f]_{W_{p,{\rm m}}^{s}({\mathbb{R}}^{d},\gamma)}<\infty\right\},$
@@
We denote by $\tilde{W}^{s}_{p,{\rm m}}(\mathbb{U}^{d})$ the subspace of $W^{s}_{p,{\rm m}}(\mathbb{U}^{d})$ consisting of functions that are periodic in each variable. The proof of the following proposition proceeds analogously to that of Proposition 3.1.

### 18 Apr 2026 v2

@@ 3 Approximation for Gaussian Sobolev spaces
$\int_{{\mathbb{I}}^{d}_{\theta}}\int_{{\mathbb{I}}^{d}_{\theta}}\frac{\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}}+{\boldsymbol{k}})-D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{y}}+{\boldsymbol{k}})\big|^{p}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+\tilde{s}p}}{\rm d}{\boldsymbol{x}}{\rm d}{\boldsymbol{y}}$
$\ll\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\int_{{\mathbb{I}}^{d}_{\theta,{\boldsymbol{k}}}}\big|D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{x}})-D^{\boldsymbol{\alpha}}f_{\boldsymbol{k}}({\boldsymbol{y}})\big|^{p}e^{\frac{|{\boldsymbol{x}}|^{2}+|{\boldsymbol{y}}|^{2}}{2}}{\rm d}\gamma({\boldsymbol{x}}){\rm d}\gamma({\boldsymbol{y}})$
@@
which together with (3.2) implies
@@
The condition $q_{*}<\frac{p_{*}}{2}$ ensures that the sum is bounded by $Cn^{-a}(\log n)^{b}\|f\|_{W^{s}_{p,G}(\mathbb{R}^{d},\gamma)}$. The remaining steps follow similarly to those in the proof for $W^{s}_{p}({\mathbb{R}}^{d},\gamma)$. The proof is complete.
@@
Proof. We present the proof for the case of Kolmogorov and linear widths; the case of sampling widths follows similarly. In this proof, for $k>0$ we denote ${\mathbb{U}}^{d}_{k}=[-k,k]^{d}$. Let $f$ be a $1$-periodic function on ${\mathbb{R}}^{d}$ and $f\in\tilde{W}^{\alpha}_{p}({\mathbb{I}}^{d})$. Denote by $f_{\operatorname{ext}}$ the extension of $f$ to $W^{s}_{p}({\mathbb{R}}^{d})$. For the existence of such a extension we refer the reader to [21, Theorem 1.105]. Without loss of generality we can assume that $\operatorname{supp}(f)\in{\mathbb{U}}^{d}_{1}$. Hence, we have
@@
Hence, we obtain
$\lambda_{n}\big({\boldsymbol{W}}^{s}_{p}(\mathbb{R}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\geq d_{n}\big({\boldsymbol{W}}^{s}_{p}(\mathbb{R}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\gg d_{n}(\tilde{\boldsymbol{W}^{s}_{p}}({\mathbb{I}}^{d}),\tilde{L}_{q}({\mathbb{I}}^{d})).$
@@
    $d_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{R}}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\asymp\lambda_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{R}}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\asymp n^{-s/d}.$
@@
    $\varrho_{n}\big({\boldsymbol{W}}^{s}_{p,G}({\mathbb{R}}^{d},\gamma),L_{q}({\mathbb{R}}^{d},\gamma)\big)\asymp n^{-s/d}.$
@@
$d_{n}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)=\lambda_{n}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big).$
@@
where the linear operators $A$, $D$, and $B$ are defined as
@@
$D\xi$ $=\big(\xi_{\boldsymbol{k}}/(1+|{\boldsymbol{k}}|_{1})^{s/2}\big)_{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}}\,,\qquad\xi=(\xi_{\boldsymbol{k}})_{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}}$
@@
$d_{n-1}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)=d_{n}(id)\leq d_{n}\big(D\big)\,.$
@@
$d_{n-1}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)=d_{n}(Id)\geq d_{n}\big(D\big)\,.$
@@
$d_{n-1}\big({\boldsymbol{H}}^{s}(\mathbb{R}^{d},\gamma),L_{2}({\mathbb{R}}^{d},\gamma)\big)=d_{n}\big(Id\big)=d_{n}\big(D\big).$
We know that $d_{n}\big(D\big)=a_{n}(D)=\sigma_{n}$ where $\sigma_{n}$ is the non-increasing rearrangement of $(1+|{\boldsymbol{k}}|_{1})^{-s/2}$. For $r\in{\mathbb{N}}$, we know that the cardinality of the set $\{{\boldsymbol{k}}\in{\mathbb{N}}^{d}_{0}:1+|{\boldsymbol{k}}|_{1}\leq r\}=$ is $c(r,d)=\binom{r-1+d}{d}$. If $c(r-1,d)<n\leq c(r,d)$. Then we have
@@
Using $c(r,d)=\binom{r-1+d}{d}$ we obtain the desired results.
@@ 4 Gaussian Sobolev spaces with mixed smoothness
For $s\in{\mathbb{N}}$, we define the Gaussian Sobolev space with mixed smoothness $W^{s}_{p,{\rm m}}({\mathbb{R}}^{d},\gamma)$ as the normed space of all functions $f\in L_{p}({\mathbb{R}}^{d},\gamma)$ such that the generalized partial derivatives $D^{\boldsymbol{\alpha}}f$ of order $\boldsymbol{\alpha}$} belong to $L_{p}({\mathbb{R}}^{d},\gamma)$ for all ${\boldsymbol{\alpha}}\in{\mathbb{N}}_{0}^{d}$ satisfying $|{\boldsymbol{\alpha}}|_{\infty}\leq s$. The norm of a function $f$ in this space is defined by
@@
Let $1\leq p<\infty$, $s>0$ and $s\not\in{\mathbb{N}}$. Then the fractional Sobolev spaces with mixed smoothness $W_{p,{\rm m}}^{s}({\mathbb{U}}^{d})$ is defined by
@@
Let $1\leq p<\infty$, $s>0$ and $s\not\in{\mathbb{N}}$. We define the fractional Gaussian Sobolev space with mixed smoothness $W_{p,{\rm m}}^{s}(\mathbb{R}^{d},\gamma)$ as
$W_{p,{\rm m}}^{s}(\mathbb{R}^{d},\gamma):=\left\{f\in W^{\bar{s}}_{p,{\rm m}}(\mathbb{R}^{d},\gamma)\;:\;[f]_{W_{p,{\rm m}}^{s}({\mathbb{R}}^{d},\gamma)}<\infty\right\},$
@@
We denote by $\tilde{W}^{s}_{p,{\rm m}}(\mathbb{U}^{d})$ the subspace of $W^{s}_{p,{\rm m}}(\mathbb{U}^{d})$ consisting of functions that are $\mathbb{U}$}-periodic in each variable. The proof of the following proposition proceeds analogously to that of Proposition 3.1.